# HOPSE: Scalable Higher-Order Positional and Structural Encoder for Combinatorial Representations

### v1

@@ F.5.1 Analysis of performance across neighborhoods
Graph Tasks Nbhd MUTAG ($\uparrow$) PROTEINS ($\uparrow$) NCI1 ($\uparrow$) NCI109 ($\uparrow$) ZINC ($\downarrow$) HOPSE-M Adj-1 88.09 $\pm$ 5.55 75.77 $\pm$ 2.67 76.85 $\pm$ 0.83 76.34 $\pm$ 1.02 0.29 $\pm$ 0.01 Adj-2 82.13 $\pm$ 4.15 74.55 $\pm$ 2.65 78.09 $\pm$ 1.44 77.27 $\pm$ 1.01 0.27 $\pm$ 0.01 Adj-3 85.53 $\pm$ 3.50 75.27 $\pm$ 2.19 77.57 $\pm$ 1.48 75.37 $\pm$ 0.79 0.28 $\pm$ 0.01 Inc-1 86.81 $\pm$ 4.61 76.99 $\pm$ 2.63 77.32 $\pm$ 1.62 75.70 $\pm$ 1.36 0.27 $\pm$ 0.02 Inc-2 83.40 $\pm$ 3.81 75.56 $\pm$ 2.26 75.89 $\pm$ 2.14 75.43 $\pm$ 1.98 0.29 $\pm$ 0.02 Inc-3 82.98 $\pm$ 3.36 74.70 $\pm$ 3.05 77.14 $\pm$ 0.46 74.46 $\pm$ 1.52 0.25 $\pm$ 0.01 Mix-1 81.28 $\pm$ 5.90 75.13 $\pm$ 3.09 77.26 $\pm$ 1.70 76.13 $\pm$ 1.38 0.25 $\pm$ 0.01 Mix-2 84.26 $\pm$ 6.13 72.62 $\pm$ 2.55 76.69 $\pm$ 1.25 76.57 $\pm$ 1.55 0.26 $\pm$ 0.01 HOPSE-G Adj-1 82.13 $\pm$ 7.76 76.27 $\pm$ 2.70 78.31 $\pm$ 0.84 77.48 $\pm$ 0.99 0.26 $\pm$ 0.00 Adj-2 87.66 $\pm$ 6.96 76.42 $\pm$ 2.19 78.83 $\pm$ 1.40 77.48 $\pm$ 0.64 0.24 $\pm$ 0.00 Adj-3 88.09 $\pm$ 1.90 75.91 $\pm$ 1.89 78.35 $\pm$ 1.11 76.94 $\pm$ 1.46 0.23 $\pm$ 0.00 Inc-1 85.11 $\pm$ 6.56 74.62 $\pm$ 3.99 77.57 $\pm$ 1.67 76.81 $\pm$ 1.23 0.25 $\pm$ 0.01 Inc-2 82.55 $\pm$ 8.82 74.91 $\pm$ 2.11 78.15 $\pm$ 0.75 76.40 $\pm$ 1.86 0.27 $\pm$ 0.01 Inc-3 84.68 $\pm$ 7.88 74.12 $\pm$ 3.39 76.81 $\pm$ 0.59 75.26 $\pm$ 1.16 0.26 $\pm$ 0.00 Mix-1 88.09 $\pm$ 3.23 76.20 $\pm$ 2.62 79.03 $\pm$ 0.73 76.88 $\pm$ 0.95 0.23 $\pm$ 0.00 Mix-2 85.11 $\pm$ 6.02 75.63 $\pm$ 2.35 79.07 $\pm$ 0.65 77.81 $\pm$ 1.26 0.21 $\pm$ 0.01
F.5.2 Analysis of time efficiency across neighborhoods
This section presents the effects of the choice of ${\mathcal{N}}_{C}$ on the total runtime and per-epoch performance of the HOPSE framework. The results in Table 8 and Table 9 report the per-epoch runtime for each neighborhood configuration across both the simplicial and cell domains, under the HOPSE-M and HOPSE-G models, respectively. While Table 10 and Table 11 present the total runtime to convergence for eight different neighbourhoods for simplicial and cell domains, and HOPSE-M and HOPSE-G correspondingly.
Table 8: Runtime per epoch per neighborhood on simplicial. Results are shown as mean $\pm$ standard deviation. The best result per domain and model is in bold and results within one standard deviation of the lowest per domain are highlighted in blue. Simpler configurations such as Adj-1 consistently yield the lowest runtime across domains, particularly for HOPSE-M, where Adj-1 achieves the fastest epoch times in 5 out of 8 datasets,
Graph Tasks Topological Tasks Nbhd MUTAG PROTEINS NCI1 NCI109 ZINC NAME ORIENT $(\beta_{0},\beta_{1},\beta_{2})$ HOPSE-M Adj-1 0.02 $\pm$ 0.00 0.12 $\pm$ 0.01 0.32 $\pm$ 0.00 0.32 $\pm$ 0.00 0.30 $\pm$ 0.01 4.12 $\pm$ 0.08 4.41 $\pm$ 0.11 4.48 $\pm$ 0.16 Adj-2 0.03 $\pm$ 0.01 0.12 $\pm$ 0.00 0.36 $\pm$ 0.01 0.36 $\pm$ 0.00 0.35 $\pm$ 0.01 4.37 $\pm$ 0.13 4.15 $\pm$ 0.09 4.54 $\pm$ 0.21 Adj-3 0.02 $\pm$ 0.00 0.12 $\pm$ 0.01 0.37 $\pm$ 0.00 0.39 $\pm$ 0.00 0.37 $\pm$ 0.01 4.42 $\pm$ 0.11 4.41 $\pm$ 0.22 4.91 $\pm$ 0.07 Inc-1 0.03 $\pm$ 0.00 0.14 $\pm$ 0.01 0.44 $\pm$ 0.01 0.40 $\pm$ 0.01 0.40 $\pm$ 0.01 6.28 $\pm$ 1.34 5.14 $\pm$ 0.24 5.41 $\pm$ 0.17 Inc-2 0.03 $\pm$ 0.01 0.12 $\pm$ 0.01 0.36 $\pm$ 0.00 0.39 $\pm$ 0.00 0.34 $\pm$ 0.01 4.41 $\pm$ 0.05 4.29 $\pm$ 0.06 4.87 $\pm$ 0.12 Inc-3 0.02 $\pm$ 0.00 0.16 $\pm$ 0.00 0.46 $\pm$ 0.01 0.47 $\pm$ 0.03 0.42 $\pm$ 0.01 5.89 $\pm$ 0.31 5.71 $\pm$ 0.23 5.70 $\pm$ 0.23 Mix-1 0.03 $\pm$ 0.00 0.17 $\pm$ 0.01 0.51 $\pm$ 0.00 0.51 $\pm$ 0.02 0.48 $\pm$ 0.01 6.00 $\pm$ 0.29 6.25 $\pm$ 0.31 6.51 $\pm$ 0.23 Mix-2 0.02 $\pm$ 0.00 0.15 $\pm$ 0.00 0.45 $\pm$ 0.01 0.44 $\pm$ 0.00 0.40 $\pm$ 0.01 5.13 $\pm$ 0.14 5.60 $\pm$ 0.55 5.54 $\pm$ 0.05 HOPSE-G Adj-1 0.02 $\pm$ 0.00 0.10 $\pm$ 0.01 0.28 $\pm$ 0.01 0.28 $\pm$ 0.00 0.26 $\pm$ 0.02 3.19 $\pm$ 0.09 3.18 $\pm$ 0.12 3.55 $\pm$ 0.12 Adj-2 0.02 $\pm$ 0.00 0.12 $\pm$ 0.01 0.30 $\pm$ 0.00 0.31 $\pm$ 0.01 0.28 $\pm$ 0.01 3.53 $\pm$ 0.12 3.41 $\pm$ 0.10 3.67 $\pm$ 0.12 Adj-3 0.02 $\pm$ 0.00 0.11 $\pm$ 0.00 0.33 $\pm$ 0.02 0.34 $\pm$ 0.01 0.29 $\pm$ 0.02 3.79 $\pm$ 0.14 3.84 $\pm$ 0.14 3.98 $\pm$ 0.07 Inc-1 0.02 $\pm$ 0.00 0.13 $\pm$ 0.01 0.38 $\pm$ 0.01 0.37 $\pm$ 0.02 0.34 $\pm$ 0.01 4.44 $\pm$ 0.14 4.23 $\pm$ 0.12 4.64 $\pm$ 0.08 Inc-2 0.02 $\pm$ 0.00 0.10 $\pm$ 0.00 0.31 $\pm$ 0.00 0.31 $\pm$ 0.00 0.28 $\pm$ 0.01 3.74 $\pm$ 0.17 3.66 $\pm$ 0.10 3.78 $\pm$ 0.06 Inc-3 0.02 $\pm$ 0.00 0.13 $\pm$ 0.00 0.41 $\pm$ 0.00 0.40 $\pm$ 0.01 0.38 $\pm$ 0.03 5.17 $\pm$ 0.40 4.75 $\pm$ 0.23 5.13 $\pm$ 0.06 Mix-1 0.02 $\pm$ 0.00 0.16 $\pm$ 0.01 0.46 $\pm$ 0.02 0.46 $\pm$ 0.01 0.44 $\pm$ 0.01 5.50 $\pm$ 0.24 5.74 $\pm$ 0.22 5.84 $\pm$ 0.38 Mix-2 0.02 $\pm$ 0.00 0.13 $\pm$ 0.01 0.38 $\pm$ 0.00 0.38 $\pm$ 0.01 0.36 $\pm$ 0.02 4.48 $\pm$ 0.14 4.55 $\pm$ 0.19 4.84 $\pm$ 0.11
Table 9: Runtime per epoch per neighborhood on cell. Results are shown as mean $\pm$ standard deviation. The best result per domain and model is in bold and results within one standard deviation of the lowest per domain are highlighted in blue.
Graph Tasks Nbhd MUTAG PROTEINS NCI1 NCI109 ZINC HOPSE-M Adj-1 0.02 $\pm$ 0.00 0.11 $\pm$ 0.00 0.31 $\pm$ 0.01 0.32 $\pm$ 0.01 0.30 $\pm$ 0.01 Adj-2 0.02 $\pm$ 0.00 0.11 $\pm$ 0.00 0.34 $\pm$ 0.00 0.33 $\pm$ 0.00 0.30 $\pm$ 0.01 Adj-3 0.02 $\pm$ 0.00 0.11 $\pm$ 0.00 0.35 $\pm$ 0.00 0.34 $\pm$ 0.00 0.32 $\pm$ 0.01 Inc-1 0.02 $\pm$ 0.00 0.13 $\pm$ 0.01 0.40 $\pm$ 0.01 0.40 $\pm$ 0.02 0.37 $\pm$ 0.02 Inc-2 0.02 $\pm$ 0.00 0.12 $\pm$ 0.00 0.35 $\pm$ 0.01 0.34 $\pm$ 0.01 0.32 $\pm$ 0.01 Inc-3 0.03 $\pm$ 0.01 0.15 $\pm$ 0.00 0.42 $\pm$ 0.01 0.43 $\pm$ 0.01 0.41 $\pm$ 0.01 Mix-1 0.03 $\pm$ 0.00 0.16 $\pm$ 0.01 0.50 $\pm$ 0.01 0.54 $\pm$ 0.01 0.47 $\pm$ 0.02 Mix-2 0.02 $\pm$ 0.00 0.13 $\pm$ 0.00 0.47 $\pm$ 0.02 0.46 $\pm$ 0.00 0.39 $\pm$ 0.01 HOPSE-G Adj-1 0.02 $\pm$ 0.00 0.09 $\pm$ 0.01 0.25 $\pm$ 0.01 0.26 $\pm$ 0.01 0.24 $\pm$ 0.00 Adj-2 0.02 $\pm$ 0.00 0.09 $\pm$ 0.00 0.29 $\pm$ 0.01 0.27 $\pm$ 0.01 0.26 $\pm$ 0.01 Adj-3 0.02 $\pm$ 0.00 0.10 $\pm$ 0.00 0.32 $\pm$ 0.01 0.32 $\pm$ 0.02 0.27 $\pm$ 0.01 Inc-1 0.02 $\pm$ 0.00 0.11 $\pm$ 0.00 0.37 $\pm$ 0.01 0.34 $\pm$ 0.00 0.33 $\pm$ 0.00 Inc-2 0.02 $\pm$ 0.00 0.09 $\pm$ 0.00 0.32 $\pm$ 0.01 0.30 $\pm$ 0.05 0.28 $\pm$ 0.02 Inc-3 0.02 $\pm$ 0.00 0.14 $\pm$ 0.01 0.40 $\pm$ 0.00 0.37 $\pm$ 0.00 0.36 $\pm$ 0.02 Mix-1 0.03 $\pm$ 0.00 0.14 $\pm$ 0.00 0.44 $\pm$ 0.01 0.45 $\pm$ 0.00 0.42 $\pm$ 0.00 Mix-2 0.02 $\pm$ 0.00 0.14 $\pm$ 0.00 0.40 $\pm$ 0.02 0.37 $\pm$ 0.01 0.32 $\pm$ 0.01
Runtime per epoch per neighborhood.
As expected Table 8 and Table 9 show that the runtime per epoch increases as the number of neighborhoods in ${\mathcal{N}}_{\mathcal{C}}$ increases. This aligns with the complexity analysis presented in Appendix E. Still, the relative increase in computational time is at most 80%, which is acceptable when considering how many more higher-order relations are being considered.
The efficiency-expressivity trade-off revealed in these results has important implications for the practical deployment. While intuition might suggest that more complex neighborhood configurations would always incur substantial computational penalties, our findings demonstrate that this is not necessarily true.
Despite incorporating significantly more structural information, complex neighborhood configurations like Mix-1 and Mix-2 introduce only a modest increase in per-epoch runtime, typically less than 2×2$\times$2 × compared to the simplest configurations. This makes them viable options even in computationally constrained environments.
Total runtime per neighborhood.
The total runtime required to reach convergence, presented in Table 10 and Table 11, reveals more nuanced patterns. Unlike per-epoch runtime, total runtime encapsulates both computational cost per epoch and the number of epochs to convergence, which are indirectly influenced by the expressiveness of the neighborhood structure. Interestingly, richer neighborhood configurations, despite their potentially higher per-epoch cost, often lead to a similar relative increase in time needed as the per-epoch case. This means that the number of epochs needed for convergence stays relatively constant.
Table 10: Total runtime per neighborhood on simplicial. Results are shown as mean $\pm$ standard deviation. The best result per domain and model is in bold and results within one standard deviation of the lowest per domain are highlighted in blue. HOPSE configurations, especially with mixed neighborhood sets, provide an excellent balance between efficiency and expressiveness.
Graph Tasks Topological Tasks Nbhd MUTAG PROTEINS NCI1 NCI109 ZINC NAME ORIENT $(\beta_{0},\beta_{1},\beta_{2})$ HOPSE-M Adj-1 6.76 $\pm$ 1.38 44.60 $\pm$ 17.17 164.10 $\pm$ 79.48 125.46 $\pm$ 19.54 1238.09 $\pm$ 378.91 2652.62 $\pm$ 67.51 3028.14 $\pm$ 41.05 5327.90 $\pm$ 1334.91 Adj-2 8.66 $\pm$ 0.51 31.26 $\pm$ 15.10 165.10 $\pm$ 31.84 131.47 $\pm$ 24.84 1512.34 $\pm$ 310.43 3036.30 $\pm$ 1105.71 2535.98 $\pm$ 45.80 4053.39 $\pm$ 718.36 Adj-3 6.87 $\pm$ 1.54 38.41 $\pm$ 8.07 175.02 $\pm$ 49.19 146.19 $\pm$ 27.88 $1$$\pm$ 404.64 2734.33 $\pm$ 156.56 2689.60 $\pm$ 38.86 4266.24 $\pm$ 651.78 Inc-1 8.52 $\pm$ 1.47 41.92 $\pm$ 5.67 194.90 $\pm$ 64.73 206.44 $\pm$ 16.44 1183.26 $\pm$ 135.82 4351.87 $\pm$ 1203.66 3342.21 $\pm$ 88.20 5172.86 $\pm$ 1084.20 Inc-2 10.37 $\pm$ 2.21 40.03 $\pm$ 20.40 161.40 $\pm$ 39.23 180.17 $\pm$ 22.35 895.12 $\pm$ 295.76 2704.44 $\pm$ 23.84 2998.32 $\pm$ 790.12 4859.76 $\pm$ 819.26 Inc-3 8.81 $\pm$ 3.00 39.99 $\pm$ 6.31 209.70 $\pm$ 38.55 218.69 $\pm$ 25.83 1332.58 $\pm$ 316.15 3493.64 $\pm$ 62.70 4076.33 $\pm$ 671.42 5465.95 $\pm$ 1816.61 Mix-1 9.81 $\pm$ 2.93 54.94 $\pm$ 13.50 214.50 $\pm$ 7.54 198.45 $\pm$ 54.21 1288.78 $\pm$ 590.16 4201.85 $\pm$ 851.36 4315.76 $\pm$ 1097.48 6542.09 $\pm$ 1272.01 Mix-2 7.63 $\pm$ 1.71 43.67 $\pm$ 18.29 229.05 $\pm$ 65.67 189.19 $\pm$ 45.39 1434.43 $\pm$ 547.60 3596.38 $\pm$ 519.37 3408.67 $\pm$ 162.45 5377.27 $\pm$ 185.49 HOPSE-G Adj-1 4.56 $\pm$ 0.62 25.49 $\pm$ 6.16 92.97 $\pm$ 27.48 104.94 $\pm$ 28.21 542.03 $\pm$ 94.68 1891.07 $\pm$ 17.81 1885.35 $\pm$ 28.06 2194.20 $\pm$ 42.70 Adj-2 4.92 $\pm$ 0.37 32.03 $\pm$ 9.91 118.32 $\pm$ 17.84 111.18 $\pm$ 18.27 480.02 $\pm$ 100.70 2172.54 $\pm$ 216.62 2096.15 $\pm$ 33.94 2323.75 $\pm$ 222.16 Adj-3 5.58 $\pm$ 1.28 39.85 $\pm$ 17.90 120.88 $\pm$ 31.47 134.65 $\pm$ 49.81 478.06 $\pm$ 107.13 2302.00 $\pm$ 47.08 2385.63 $\pm$ 176.04 2997.61 $\pm$ 824.53 Inc-1 5.80 $\pm$ 0.27 32.71 $\pm$ 8.43 130.06 $\pm$ 22.37 106.98 $\pm$ 27.98 649.05 $\pm$ 114.28 2580.45 $\pm$ 47.63 2471.89 $\pm$ 47.46 3716.82 $\pm$ 1192.83 Inc-2 6.44 $\pm$ 0.54 32.98 $\pm$ 8.25 128.01 $\pm$ 22.39 99.39 $\pm$ 33.07 661.57 $\pm$ 133.52 2489.63 $\pm$ 444.47 2240.50 $\pm$ 43.84 2568.75 $\pm$ 76.80 Inc-3 6.85 $\pm$ 0.98 51.18 $\pm$ 11.23 174.20 $\pm$ 57.02 148.67 $\pm$ 37.09 778.78 $\pm$ 117.23 3038.25 $\pm$ 65.40 2930.43 $\pm$ 430.59 3602.56 $\pm$ 555.32 Mix-1 6.85 $\pm$ 0.72 51.94 $\pm$ 9.59 209.07 $\pm$ 28.40 150.07 $\pm$ 36.64 911.69 $\pm$ 165.22 3408.18 $\pm$ 515.16 3379.79 $\pm$ 111.99 4321.64 $\pm$ 1094.92 Mix-2 5.16 $\pm$ 0.36 35.97 $\pm$ 4.38 168.30 $\pm$ 38.75 142.36 $\pm$ 34.54 716.46 $\pm$ 53.19 3324.16 $\pm$ 1113.48 2794.73 $\pm$ 345.73 3876.34 $\pm$ 1141.15
@@ F.5.2 Analysis of time efficiency across neighborhoods
Table 11: Total runtime per neighborhood on cell. Results are shown as mean $\pm$ standard deviation. The best result per domain and model is in bold and results within one standard deviation of the lowest per domain are highlighted in blue. On HOPSE-G, some more complex neighborhoods will result in faster time, even though they contain more information and thus parameters to be learned. This is one of the effects of convergence speed. This can be seen in neighborhoods Mix-1 and Mix-2. On HOPSE-M, there is no such drastic difference
Graph Tasks Nbhd MUTAG PROTEINS NCI1 NCI109 ZINC HOPSE-M Adj-1 7.03 $\pm$ 0.86 37.52 $\pm$ 2.06 144.11 $\pm$ 32.07 142.87 $\pm$ 45.01 1128.14 $\pm$ 166.57 Adj-2 7.66 $\pm$ 0.82 35.68 $\pm$ 7.88 147.04 $\pm$ 24.07 121.89 $\pm$ 18.30 901.44 $\pm$ 203.17 Adj-3 9.78 $\pm$ 2.94 37.85 $\pm$ 10.19 153.94 $\pm$ 29.39 131.39 $\pm$ 36.31 1414.14 $\pm$ 325.91 Inc-1 7.86 $\pm$ 0.54 37.06 $\pm$ 7.05 174.19 $\pm$ 33.02 190.50 $\pm$ 35.45 1173.43 $\pm$ 198.41 Inc-2 6.41 $\pm$ 1.43 36.82 $\pm$ 11.82 220.47 $\pm$ 55.22 128.32 $\pm$ 14.86 1020.04 $\pm$ 131.15 Inc-3 8.30 $\pm$ 0.78 48.24 $\pm$ 18.26 169.79 $\pm$ 41.60 169.78 $\pm$ 37.03 1183.44 $\pm$ 217.72 Mix-1 8.83 $\pm$ 1.18 51.25 $\pm$ 13.72 205.54 $\pm$ 25.51 210.50 $\pm$ 11.62 1313.43 $\pm$ 132.16 Mix-2 7.42 $\pm$ 1.37 44.34 $\pm$ 13.40 203.46 $\pm$ 48.75 193.17 $\pm$ 44.83 1676.92 $\pm$ 398.90 HOPSE-G Adj-1 5.09 $\pm$ 0.93 26.71 $\pm$ 4.79 88.31 $\pm$ 11.27 95.95 $\pm$ 28.16 683.86 $\pm$ 125.70 Adj-2 4.98 $\pm$ 0.47 31.85 $\pm$ 5.43 96.33 $\pm$ 19.91 93.98 $\pm$ 15.00 887.83 $\pm$ 160.74 Adj-3 6.64 $\pm$ 0.89 31.13 $\pm$ 5.56 100.65 $\pm$ 14.14 113.39 $\pm$ 18.75 595.39 $\pm$ 107.46 Inc-1 5.75 $\pm$ 1.45 34.11 $\pm$ 7.89 110.19 $\pm$ 19.33 105.67 $\pm$ 27.00 767.42 $\pm$ 85.39 Inc-2 7.65 $\pm$ 2.32 24.65 $\pm$ 6.74 105.02 $\pm$ 28.61 126.14 $\pm$ 37.05 721.57 $\pm$ 60.10 Inc-3 7.87 $\pm$ 1.47 44.42 $\pm$ 9.81 181.18 $\pm$ 96.45 100.22 $\pm$ 11.50 896.86 $\pm$ 162.01 Mix-1 7.61 $\pm$ 0.68 46.14 $\pm$ 10.02 239.59 $\pm$ 38.24 168.35 $\pm$ 50.01 671.40 $\pm$ 948.29 Mix-2 5.27 $\pm$ 0.57 36.29 $\pm$ 13.05 152.73 $\pm$ 8.26 104.12 $\pm$ 37.76 1229.56 $\pm$ 527.27
Optimal configurations.
For simpler graph classification tasks, simpler neighborhood configurations may offer the best efficiency. In contrast, more expressive neighborhood sets often provide better efficiency-accuracy trade-offs for complex topological tasks and larger datasets. These findings demonstrate that the choice of neighborhood configuration represents a critical design decision not just for model accuracy, but also for computational efficiency. The provided analysis presents a principled approach to neighborhood selection based on task requirements and domain characteristics, establishing an essential foundation for applying TDL to new fields.
F.5.3 Preprocessing time per neighborhood
Preprocessing time analysis
The computational costs of extracting positional and structural encodings are summarized in Tables 12 and 13. While preprocessing times scale with the complexity of the selected neighborhoods and not only with the number of cells, two points need to be emphasized:
Table 12: Preprocessing time per neighborhood on simplicial. Results are shown as mean time in seconds. The best result per domain and model is in bold. All Topological Tasks have the same time as they share the same preprocessing time. Times increase with neighborhood since with Adj-1 being the fastest as it’s the neighborhood of the underlying graph.

### v2

@@ Appendix H Performance ablation across different neighborhoods
Graph Tasks $\mathcal{N}_{S}$ MUTAG ($\uparrow$) PROTEINS ($\uparrow$) NCI1 ($\uparrow$) NCI109 ($\uparrow$) ZINC ($\downarrow$) HOPSE-M Adj-1 $83.40\pm 0.04$.85) $75.05\pm 0.02$.33) $76.09\pm 0.01$.99) $76.38\pm 0.00$.73) $0.27\pm 0.00$.01) Adj-2 $83.83\pm 0.03$.87) $77.28\pm 0.02$.44) $76.71\pm 0.00$.42) $75.80\pm 0.01$.71) $0.26\pm 0.00$.01) Adj-3 $85.53\pm 0.04$.85) $76.06\pm 0.03$.24) $76.05\pm 0.01$.75) $75.68\pm 0.00$.93) $25\pm 0.00.00$) Inc-1 $83.83\pm 0.05$.12) $76.99\pm 0.01$.94) $77.88\pm 0.01$.71) $76.32\pm 0.01$.40) $0.26\pm 0.00$.01) Inc-2 $86.38\pm 0.05$.34) $75.84\pm 0.02$.94) $76.63\pm 0.01$.78) $76.03\pm 0.01$.45) $0.26\pm 0.00$.01) Inc-3 $80.43\pm 0.05$.51) $75.70\pm 0.03$.55) $76.77\pm 0.00$.21) $75.78\pm 0.01$.66) $0.25\pm 0.00$.02) Mix-1 $84.26\pm 0.06$.49) $76.63\pm 0.01$.99) $77.80\pm 0.00$.82) $77.08\pm 0.01$.35) $0.24\pm 0.00$.01) Mix-2 $83.83\pm 0.04$.15) $77.06\pm 0.02$.43) $76.52\pm 0.02$.26) $76.22\pm 0.01$.64) $0.24\pm 0.00$.01) HOPSE-G Adj-1 $82.13\pm 0.07$.76) $76.27\pm 0.02$.70) $78.31\pm 0.00$.84) $77.48\pm 0.00$.99) $26\pm 0.00.00$) Adj-2 $87.66\pm 0.06$.96) $76.42\pm 0.02$.19) $78.83\pm 0.01$.40) $77.48\pm 0.00$.64) $24\pm 0.00.00$) Adj-3 $88.09\pm 0.01$.90) $75.91\pm 0.01$.89) $78.35\pm 0.01$.11) $76.94\pm 0.01$.46) $23\pm 0.00.00$) Inc-1 $85.11\pm 0.06$.56) $74.62\pm 0.03$.99) $77.57\pm 0.01$.67) $76.81\pm 0.01$.23) $0.25\pm 0.00$.01) Inc-2 $82.55\pm 0.08$.82) $74.91\pm 0.02$.11) $78.15\pm 0.00$.75) $76.40\pm 0.01$.86) $0.27\pm 0.00$.01) Inc-3 $4.68\pm 0.07.88$) $74.12\pm 0.03$.39) $76.81\pm 0.00$.59) $75.26\pm 0.01$.16) $26\pm 0.00.00$) Mix-1 $88.09\pm 0.03$.23) $76.20\pm 0.02$.62) $79.03\pm 0.00$.73) $76.88\pm 0.00$.95) $23\pm 0.00.00$) Mix-2 $85.11\pm 0.06$.02) $75.63\pm 0.02$.35) $79.07\pm 0.00$.65) $77.81\pm 0.01$.26) $0.21\pm 0.00$.01)
Appendix I Runtime ablation across different neighborhoods
I.1 Total runtime
@@
On HOPSE-G, some more complex neighborhoods will result in faster time, even though they contain more information and thus parameters to be learned. This is one of the effects of convergence speed. This can be seen in neighborhoods Mix-1 and Inc-[1,3]. On HOPSE-M, there is no such drastic difference in HOPSE configurations, especially with mixed neighborhood sets, which provide an excellent balance between efficiency and expressiveness.
Table 8: Total training runtime (seconds) ablation of $\mathcal{N}_{S}$ on simplicial domain.